### v1

@@ Abstract
Deploying a Machine Learning (ML) training pipeline into production requires robust software engineering practices. This differs significantly from experimental workflows. This experience report investigates this challenge in SPIRA, a project whose goal is to create an ML-Enabled System (MLES) to pre-diagnose insufficiency respiratory via speech analysis. The first version of SPIRA’s training pipeline lacked critical software quality attributes. This paper presents an overview of the MLES, then compares three versions of the architecture of the Continuous Training subsystem, which evolved from a Big Ball of Mud, to a Modular Monolith, towards Microservices. By adopting different design principles and patterns to enhance its maintainability, robustness, and extensibility. In this way, the paper seeks to offer insights for both ML Engineers tasked to productionize ML training pipelines and Data Scientists seeking to adopt MLOps practices.
@@
Code Quality MLOps Software Architecture Machine Learning Enabled Systems Healthcare Domain Case Study
@@ 1 Introduction
Since then, the scope of SPIRA has evolved: its proposal is to detect respiratory insufficiency of different origins. Multiple sickness can cause this symptom, including smoking side effects, flu, severe asthma, and heart conditions [5].
@@
The paper is divided as follows. Section 2 presents the architecture of the MLES. Section 3 refers to previous research and narrows the scope to incremental development of the Continuous Training subsystem. Sections 4 to 6 explain how its architecture evolved from a Big Ball of Mud [6] to a Modular Monolith [15] to Microservices [14]. Sections 7 and 8 lists challenges and lessons learned. Finally, section 9 summarizes contributions and future work.
@@ 2 The SPIRA ML-Enabled System
Figure 1: Architecture of the SPIRA ML-Enabled System. This architecture follows the same notation presented in Ferreira et al. [3]. Rectangles represent services, which execute continuously. Stacked rectangles represent pipelines, which execute a task on demand. Lastly, cylinders represent data storage, which may be databases of any type. Components are connected by arrows. Black arrows with a hollow tip illustrate the execution flow. They start and end in a component. Labeled arrows represent the trigger that starts a workflow, whereas unlabeled arrows represent the continuation of an existing workflow. Colored arrows with a filled tip illustrate the data flow. They appear in two types: solid arrows going to and from a data storage represent write and read operations, respectively; dotted arrows represent a sync or async request-response communication between components. Components are colored according to the data they produce: raw data  , ML-specific data  , source code  , executable artifacts  , ML models  , ML training metadata  , ML model predictions  , and ML model metrics  . Remaining standalone components  orchestrate the execution of others. Components are also grouped into subsystems, with their background colored according to the students responsible for their development. Numbers  , roman numerals  and letters  are used as labels throughout section 2.
@@ 2.2 Continuous Training
Casanova et al.proposed a Deep Learning (DL) architecture based on Convolutional Neural Networks (CNNs) to detect respiratory insufficiency from audio signals [2]. As more data becomes available as ground truth, there is potential to retrain the model and improve its accuracy [9].
@@ 2.6 Monitoring
Figure 2: Incremental Development of the SPIRA Continuous Training Subsystem. The first version was published by Casanova et al. [2]. It has two main scripts: one to create datasets, and one to train the CNN model. Its code has an unplanned, organically developed structure, as it often occurs in experimental code developed by data scientists  [10, 16]. Therefore, the Experimental Architecture (v1) can be classified as a Big Ball of Mud [6]. The second version was created in the bachelor thesis of Daniel Lawand. It applies multiple design patterns [7] such as Chain of Responsability, Strategy, and Template Method to decouple different responsibilities within the pipeline. Moreover, the application logic was decoupled from external dependencies with the Ports and Adapters pattern [11]. Therefore, the Modularized Architecture (v2) can be classified as a Modular Monolith [15]. The third version was created in the bachelor thesis of Lucas Quaresma and Roberto Bolgheroni. It further expands on the second version by splitting the single application into multiple to facilitate the resilient execution of the pipeline in association with a Scheduler Service (9). Therefore, the Productionized Architecture (v3) follows a Microservices architectural style [14].
@@ 3 Incremental Implementation
The SPIRA MLES described in section 2 has been incrementally developed since 2021 via scientific initiation projects and bachelor theses supervised by M.Sc. Renato Cordeiro Ferreira and Prof. Dr. Alfredo Goldman at the University of São Paulo (USP). The scope of these projects is illustrated in figure 1.
In 2022, our team published the paper “SPIRA: Building an Intelligent System for Respiratory Insufficiency Detection” [4]. That experience report summarized challenges and lessons learned of developing the Data Collection and Serving subsystems, with contributions of the students Francisco Wernke and Vitor Tamae. This paper focuses on the Continuous Training subsystem, whose development process is summarized in figure 2.
@@ 4 Experimental Architecture (v1)
The first version of the Continuous Training pipeline was created as a proof-of-concept to showcase the SPIRA model. The code was published as part of the reproduction package [^2] https://github.com/Edresson/SPIRA-ACL2021 by Casanova et al.in 2020 [2].
The package has two main sets of scripts: one to create datasets (available in the scripts folder), and one to train the CNN model (available in the models folder). The code presents an unplanned, organically developed structure, as it often occurs in the experimentation phase by data scientists [10, 16]. Figure 3 illustrates an example of the Misplaced Responsability bad smell in the code, evidence of high coupling and low cohesion in the code. Therefore, this architecture can be classified as a Big Ball of Mud [6].
@@
[ numbers=both, numbersep=0.5em, fontsize=, frame=lines, framesep=0.5em, highlightcolor=Violet!10, labelposition=topline, highlightlines= 2-7, 8-16, 18-20, 24-36, 32 , ]pythoncode/misplaced_responsability.py
Figure 3: Example of the Misplaced Responsibility bad smell. This code snippet from the Experimental Architecture (v1) shows part of the constructor of the auxiliary class Dataset. Lines 2-7 handle random number generation. Lines 8-16 assign values to attributes. Lines 18-20 and 24-27 handle assertions. Line 32 handles data loading from a CSV. The high coupling and low cohesion makes it difficult to create new experiments.
@@ 5 Modularized Architecture (v2)
The second version of the Continuous Training pipeline was created during the bachelor thesis research of Daniel Lawand, co-author of this paper. The v2 code is available open source at GitHub [^3] https://github.com/spirabr/SPIRA-training-prototype.
The system applies multiple design patterns [7] to increase the flexibility of the code. Some examples include: Chain of Responsability, to allow to dynamically structure the steps of the pipeline depending on the experiment that will be executed; Strategy, to allow choosing different data preprocessing, feature engineering, and model evaluation techniques; and Template Method, to allow the reuse of generic code snippets that can be extended for future experiments (often in combination with the Strategy pattern).
In another level of abstraction, the business logic was decoupled from external dependencies by using the Ports and Adapters architectural pattern [11], using Dependency Injection to connect different layers of the application. Therefore, the v2 architecture can be classified as a Modular Monolith [15].
@@ 6 Productionized Architecture (v3)
The third version of the Continuous Training pipeline was created during the bachelor thesis research of Lucas Quaresma and Roberto Bolgheroni, co-authors of this paper. The v3 code is available open source at GitHub [^4] https://github.com/spirabr/SPIRA-training.
The system expands on the Modularized Architecture (v2) by splitting the single application into multiple applications to facilitate the resilient execution of the pipeline in association with a Scheduler Service (9). If a failure occurs in one stage of the pipeline, it does not require re-executing it from start. Therefore, the v3 architecture follows a Microservices architectural style [14].
Testability was a key priority in this version. Following testing practices for microservices [15], the business logic received unit test, developed using Test-Driven Development (TDD), while external dependencies received integration tests, developed around the Ports and Adapters [11].
@@ 7 Challenges
The creation of the Experimental Architecture (v1) followed a typical data science workflow. Their goal is to create a proof of concept that demonstrates a problem can be solved using ML techniques [12]. In these workflows, serving a model, i.e., going into production, happens only on the last step of the process.
The Machine Learning Engineer (MLE) emerged as a specialization of a software engineer whose responsibility if to help to productionize such a model, helping the data scientist to think about the robustness of their model, and how it will fit an ML-Enabled System as a whole [1, 16].
Within this context, there were two main challenges during the evolution of the SPIRA Continuous Training architecture.
@@
The general lack of cohesion in the v1 architecture made it extremely difficult to reuse its code. The process to create the v2 architecture became an exercise of Software Archeology [8], in which the code had to be read line by line to decipher intentions, since there were no clear abstractions and some responsibilities were spread throughout the code.
Testability.
Even after the introduction of patterns in the v2 architecture, in particular the Ports and Adapters pattern, testing was still challenging. With many intents represented in the original v1 architecture, it was difficult to recover what behavior should be tested for long-term maintainability once the v3 architecture was implemented.
@@ 8 Lessons Learned
The development of ML systems is inherently difficult, and developing them without planning makes their complexity even greater. There were two main lessons learned during the process of migrating the architecture.
Collaboration of Data Scientists and ML Engineers.
By bringing these two roles together since the beginning of the experimentation process, it is easier to design a more structured code that can be maintained long term.
Testability as a first-class concern.
If the code is planned with tests and validation since its inception, it is easier to create new experiments in different components of the pipeline.
@@ 9 Conclusion
This paper reported the experience of incrementally developing the Continuous Training subsystem of SPIRA, a project that aims to develop a MLES to detect respiratory insufficiency via speech analysis. Its architecture evolved in three stages: from a proof of concept Big Ball of Mud to an organized Modular Monolith to a tested set of Microservices. By learning from this process, other projects may avoid similar pitfalls to achieve production-readiness sooner.

### v2

@@ Abstract
Deploying a Machine Learning (ML) training pipeline into production requires good software engineering practices. Unfortunately, the typical data science workflow often leads to code that lacks critical software quality attributes. This experience report investigates this problem in SPIRA, a project whose goal is to create an ML-Enabled System (MLES) to pre-diagnose insufficiency respiratory via speech analysis. This paper presents an overview of the architecture of the MLES, then compares three versions of its Continuous Training subsystem: from a proof of concept Big Ball of Mud (v1), to a design pattern-based Modular Monolith (v2), to a test-driven set of Microservices (v3) Each version improved its overall extensibility, maintainability, robustness, and resiliency. The paper shares challenges and lessons learned in this process, offering insights for researchers and practitioners seeking to productionize their pipelines.
@@
Code Quality MLOps Software Architecture Machine Learning Enabled Systems Healthcare Domain Experience Report
@@ 1 Introduction
Since then, the scope of the SPIRA project has evolved to detect respiratory insufficiency of different origins, including many sicknesses that can cause this symptom: smoking side effects, flu, severe asthma, and heart conditions [7].
@@ 2 The SPIRA ML-Enabled System
Figure 1: Architecture of the SPIRA ML-Enabled System. The architecture is described with the same notation of the reference architecture described in Ferreira et al. [5]. Rectangles represent services, which execute continuously. Stacked rectangles represent pipelines, which execute a task on demand. Lastly, cylinders represent data storage, which may be databases of any type. Components are connected by arrows. Black arrows with a hollow tip illustrate the execution flow. They start and end in a component. Labeled arrows represent the trigger that starts a workflow, whereas unlabeled arrows represent the continuation of an existing workflow. Colored arrows with a filled tip illustrate the data flow. They appear in two types: solid arrows going to and from a data storage represent write and read operations, respectively; dotted arrows represent a sync or async request-response communication between components. Components are colored according to the data they produce: raw data  , source code  , executable artifacts  , ML-specific data  , ML models  , ML training metadata  , ML model predictions  , and ML model metrics  . The remaining standalone components  orchestrate the execution of others. Components are also grouped into subsystems, with their background colored according to the students responsible for their development. Numbers  , roman numerals  and letters  are used as labels throughout section 2.
@@ 2.2 Continuous Training
Casanova et al. proposed a Deep Learning (DL) architecture based on a Convolutional Neural Network (CNN) to detect respiratory insufficiency from audio signals [4]. As more data becomes available as ground truth, there is potential to retrain the model and improve its accuracy [10].
@@ 3 Incremental Implementation
The SPIRA MLES has been incrementally developed since 2021. This paper focuses on the Continuous Training subsystem, whose development process is summarized in figure 2. For challenges and lessons learned while developing the Data Collection and Serving, subsystems, please refer to our paper “SPIRA: Building an Intelligent System for Respiratory Insufficiency Detection” [6].
Figure 2: Incremental Development of the SPIRA Continuous Training Subsystem. Dashed boxes  represent the versions of the subsystem. Rounded boxes represent deployment units. Squared boxes represent components. Black arrows with a filled tip illustrate the data flow, while the labels in italic represent the data types.
@@ 4 Experimental Architecture (v1)
The first version of the Continuous Training pipeline was created as a proof-of-concept to showcase the SPIRA model. The code was published as part of the reproduction package by Casanova et al. in 2020 [4].
The package has two main sets of scripts: one to create datasets (available in the scripts folder), and one to train the CNN model (available in the models folder). The code presents an unplanned, organically developed structure, as it often occurs in the experimentation phase by data scientists [12, 21].
Figure 3 (left) illustrates an example of the Misplaced Responsibility bad smell [15] in the code, evidence of high coupling and low cohesion in the code. In this way, this architecture can be classified as a Big Ball of Mud [8].
@@
[ label=( v1 architecture ), highlightlines= 1-6, 8-16, 18-19, 22-24, 30 , ]code/misplaced_responsability.py
[…]
[ label=( v2 architecture ), highlightlines= 3-7, 15-23, 27, 29-36 , ]code/pipeline.py
Figure 3: Example of improving modularization between v1 and v2 architectures. The v1 snippet (left) shows a Misplaced Responsibility bad smell [15] at the class Dataset. Lines 1-7 handle random number generation. Lines 8-16 assign values to attributes. Lines 18-19 and 22-24 handle assertions. Line 30 handles data loading. The v2 snippet (right) shows the application of design patterns at the module pipeline. Line 15-23 handles data loading using the Audio Adapter. Line 27 builds an audio_processor via a Chain of Responsibility pattern. Line 29-36 build feature_transformers via the Strategy pattern. Improving the modularization makes the intention of the code more explicit: allow multiple experiments depending on the configuration.
@@ 5 Modularized Architecture (v2)
The second version of the Continuous Training pipeline was created during the bachelor thesis of Daniel Lawand [14]. Its code is open source at GitHub.
This version applied multiple design patterns [9] to modularize the code. Some examples include: Chain of Responsibility, to dynamically choose the pipeline steps to be executed; Strategy, to dynamically choose techniques for data preprocessing, feature engineering, and model evaluation; and Template Method, to reuse generic code snippets. In another level of abstraction, the business logic was decoupled from external dependencies by using the Ports and Adapters architectural pattern [16], using Dependency Injection to connect different layers of the application.
Figure 3 (right) illustrates the use of design patterns. This reimplementation made the code more maintainable and extensible [19]. The v2 architecture can be classified as a Modular Monolith [20].
@@ 6 Productionized Architecture (v3)
The third version of the Continuous Training pipeline was created during the bachelor thesis of Lucas Quaresma and Roberto Bolgheroni [13]. Its code is open source at GitHub.
This version expands on the v2 architecture by splitting the single monolithic application into multiple deployment units, as described by the Workflow Pipeline ML design pattern [12]. The goal was to enable executing the pipeline with a Scheduler Service (9). In this way, if a failure occurs in one stage of the pipeline, it does not require re-executing it from the start, a costly operation.
Testability was also a key priority in this version. It was only possible because the v2 architecture made clearer the expected behaviors from the v1 architecture. Following testing practices for microservices [20], the business logic received unit tests via Test-Driven Development (TDD) [1], while the Ports and Adapters received integration tests around external dependencies [16].
This reimplementation made the code more robust and resilient [19]. The v3 architecture follows a Microservices architectural style [18].
@@ 7 Challenges
The Experimental Architecture (v1) was built following a typical CRISP-DM process [3, 21]. The goal was to create a proof-of-concept model. Serving it, i.e., going into production, was the last step of the workflow.
This development process brought a series of drawbacks for the Continuous Training subsystem’s architectural characteristics, in particular its extensibility, maintainability, robustness, and resiliency. As a consequence, there were two main challenges to migrating the subsystem toward production.
@@
The v1 architecture had many examples of Misplaced Responsibility bad smell [15], which affected its overall maintainability and extensibility [19]. Building the v2 architecture became an exercise in Software Archeology [11]: it was reimplemented from the v1 code by carefully reading it line by line to decipher its intentions, applying design patterns to decrease coupling and increase cohesion in its abstractions.
Automated Testing.
The v2 architecture had improved modularity, but it still lacked robustness and resiliency [19]. Building the v3 architecture became an exercise in Test-Driven Development (TDD) [1]: it was reimplemented from v2 code by carefully interpreting intentions around its design, creating automated tests to document behaviors that should be maintained for the long term.
@@ 8 Lessons Learned
The development of MLES is inherently difficult, and developing them without adequate planning makes their complexity even greater. As a consequence, there were two lessons learned that could have improved the migration of the Continuous Training subsystems toward production.
Collaboration between Data Scientists and ML Engineers.
Machine Learning Engineering (MLE) is a new subarea of software engineering whose goal if to help to productionize ML models [3, 21]. This discipline helps the data scientist to think about how it will fit into an MLES. By bringing these two roles together since the beginning, it is easier to make code maintainable and extensible.
Testing as a First-Class Concern.
Creating automated tests is recognized as a good practice in software engineering [2]. This discipline helps developers to think about the long-term maintenance of a system. By designing tests and validation since the beginning, it is easier to make the code robust and resilient.
@@ 9 Conclusion
This paper addressed the incremental development of SPIRA, focusing on its Continuous Training subsystem. Its architecture evolved in three stages: from a proof of concept Big Ball of Mud (v1), to a design pattern-based Modular Monolith (v2), to a test-driven set of Microservices (v3). Each step helped the subsystem’s extensibility, maintainability, robustness, and resiliency.
By learning from this experience, similar projects may employ the above lessons learned to avoid similar challenges, thus reaching production sooner,
Acknowledgements.
This research was funded by FAPESP (São Paulo Research Foundation), grant number 2023/00488-5.